chuckthedonut:

Solve for x. 2 1/2x−3/4(2x+5)=3/8 Answer choices: A) x=−4 5/8 B) x=−3 3/8 C) x=4 1/8 D) x=5 3/8

chuckthedonut:

chuckthedonut:

i literally suck at this and this is the only question i need help with, so any help would be nice :)

chuckthedonut:

okay, nevermind i'll try and figure it out myself.

Luigi0210:

Did you figure it out? Lemme answer it for you anyways, so we start off with this: \[\large 2\frac{1}{2}x -\frac{3}{4}(2x+5)=\frac{3}{8} \] First, you have to distribute the 3/4 out into the parentheses: \[\large 2\frac{1}{2}x -(\frac{3}{4}(2x)+\frac{3}{4}(5))=\frac{3}{8} \] Don't forget that it's *negative* 3/4 so: \[\large 2\frac{1}{2}x -\frac{3}{4}(2x)-\frac{3}{4}(5)=\frac{3}{8} \] Now just multiply it out: \[\large 2\frac{1}{2}x -\frac{6}{4}x-\frac{15}{4}=\frac{3}{8} \] The 6/4 can be simplified, and the 2(1/2) can be changed: \[\large \frac{5}{2}x -\frac{3}{2}x-\frac{15}{4}=\frac{3}{8} \] Now, just move -15/4 to the other side of the equation, and combine the x terms: \[\large \frac{2}{2}x=\frac{3}{8}+\frac{15}{4} \] multiply the 15/4 by 2 and you can add the fractions; and 2/2 is just one, so you get: \[\large x=\frac{33}{8} \] Since 8 can go into 33, four times, you can pull out a 4: \[\large x=4\frac{1}{8} \]

chuckthedonut:

@luigi0210 wrote:
Did you figure it out? Lemme answer it for you anyways, so we start off with this: \[\large 2\frac{1}{2}x -\frac{3}{4}(2x+5)=\frac{3}{8} \] First, you have to distribute the 3/4 out into the parentheses: \[\large 2\frac{1}{2}x -(\frac{3}{4}(2x)+\frac{3}{4}(5))=\frac{3}{8} \] Don't forget that it's *negative* 3/4 so: \[\large 2\frac{1}{2}x -\frac{3}{4}(2x)-\frac{3}{4}(5)=\frac{3}{8} \] Now just multiply it out: \[\large 2\frac{1}{2}x -\frac{6}{4}x-\frac{15}{4}=\frac{3}{8} \] The 6/4 can be simplified, and the 2(1/2) can be changed: \[\large \frac{5}{2}x -\frac{3}{2}x-\frac{15}{4}=\frac{3}{8} \] Now, just move -15/4 to the other side of the equation, and combine the x terms: \[\large \frac{2}{2}x=\frac{3}{8}+\frac{15}{4} \] multiply the 15/4 by 2 and you can add the fractions; and 2/2 is just one, so you get: \[\large x=\frac{33}{8} \] Since 8 can go into 33, four times, you can pull out a 4: \[\large x=4\frac{1}{8} \]
this was the correct answer, i got 5/5 on it ! thank you ! :D